# A Large-Scale Privacy Assessment of Android Third-Party SDKs

### v1

@@ 4.2. Collecting Android SDKs
We recognize the Google Play SDK Index as the main repository to collect existing Android SDKs. Moreover, we notice the uniqueness of mainland China’s Android ecosystem due to the absence of Google Play infrastructure, and therefore, we resort to CAICT as a supplement source to collect Android SDKs. We remark that, at the moment of this study being conducted, both sources serve as indexes that include SDKs meta information and download links but do not enforce a strict regulation and code auditing mechanism like Google Play App Store. Therefore, there is no guarantee that each SDK included in these two sources comes with comprehensive documents and a complete privacy policy requested by relevant regulations such as GDPR. In CAICT’s repository, each release version of an SDK is published separately. We need to filter out the historical versions of the duplicated SDKs and only download their latest release. As a result, we managed to download 158 SDKs as of October 2022, among which 139 are from the Google Play SDK Index and 19 are released on the CAICT website. These 158 constitute our previously defined $\mathbb{S}$ (Section 2.2) for the later assessment. We also revisit the latest versions of the same group of SDKs in October 2023 and conduct an additional assessment, which will be detailed in Section 7.
@@
Next, we implement a crawler with Selenium (selenium, ) to retrieve meta information and privacy policy links from the two sources. The meta information includes the SDK name, provider (developer’s identity), Maven IDs, and functional categories. Such information will be jointly used to uniquely identify SDKs during our subsequent analysis. We then collect the privacy policies from the retrieved links. We will use them to extract the list of personal data that the SDK requests.
The retrieval of privacy policies is not as straightforward as what we do for meta information because there does not exist a compulsory standard for developers to draft their privacy policies. Thus, for the SDKs from the Google Play SDK Index, we resort to the link indexed by data safety section guidance, if provided by developers, and download the linked webpage. Similarly, we download the webpage from the URL link indexed by SDK privacy policy link [^1] We resort to Google Webpage Translate to process all non-English text from the involved webpages. in the SDK details webpage for those SDKs collected from the CAICT website. Due to the low availability of privacy policies, we only managed to collect the privacy policies of 52 SDKs directly from the two platforms. For the remaining SDKs that do not provide a URL link to their privacy policies, we manually search the relevant documents on the Internet with a few keywords including “disclaimer”, “policy”, “disclosure”, and “guidance”. If the documents have been successfully found, we replenish the webpage into our privacy policy dataset. As a result, we managed to acquire valid privacy policy webpages for 109 SDKs in total. The text of privacy policies is extracted for our next step analysis.
4.3. Identifying Data Requested in Privacy Policies
After collecting the 109 privacy policies, our next step aims to identify the data requested by each SDK from them. In contrast to traditional app privacy policies that are written in user-oriented language, SDK privacy policies tend to be developer-oriented and may contain various terminology that are domain-specific (e.g., ICCID, OAID, BSSID) or technically synonymous (e.g., Google Ad ID, which also appears as Advertising ID, or GAID). This characteristic makes it challenging for existing methods that build privacy policy ontologies using trained models (e.g., Name Entity Recognition) (andow2019policylint, ; andow2020actions, ; bui2021consistency, ) and analyzing privacy policy frameworks based on specific sentence patterns (xie2022scrutinizing, ; yu2016can, ; zimmeck2016automated, ) to extract domain-specific entities accurately. Therefore, we leverage a LLM to address the technical challenges posed by existing approaches, and we resort to Natural Language Inference (NLI) techniques to infer the data collection from the privacy policy documents.
Figure 2. The workflow of our privacy policy analysis to find out what an SDK claims to collect
@@ 4.3. Identifying Data Requested in Privacy Policies
NLI has shown promising capability in understanding the context of natural language in literature of other domains (harkous2022hark, ). In this work, we leverage the roberta-NLI model, [^2] The full name of the model is roberta-large-snli_mnli_fever_anli _R1_R2_R3-nli. We shortly write as roberta-NLI to save the space. a state-of-the-art transformer-based LLM specialized in NLI tasks (nie2020adversarial, ), to help us identify data requested from the collected privacy policies. We briefly illustrate our approach in Figure 2.
In the usage of the roberta-NLI model, an NLI task aims to determine whether one text $\mathcal{A}$, also known as premise, could imply or infer another text $\mathcal{B}$, also known as hypothesis. By providing both texts to the NLI model, it outputs predictions of three types of relationships, as detailed below:
  • Entailment: $\mathcal{A}$ can logically imply $\mathcal{B}$.
  • Contradiction: there is a clear logical contradiction between $\mathcal{A}$ and $\mathcal{B}$.
  • Neutral: $\mathcal{A}$ and $\mathcal{B}$ do not exhibit a clear entailment or contradiction relationship.
Inference Setups
We first manually formulate hypotheses based on the 41 pre-identified data types (refer to Section 3). After that, we parse all sentences from the SDK privacy policies, employ them as the premises, and pass them to the NLI model to determine the entailment relationship with the hypotheses.
The formulation of hypotheses is shown as a non-trivial process due to two challenges found during our study. The first challenge is the developer-oriented writing style. For that reason, given a data type, we have to provide as many detailed items that fall into the same scope as possible to form the object of a hypothesis sentence. For example, both the user name and email are compositions of the account information of an Android device. Moreover, we manually collect all possible technically synonymous terms for each data type to pursue satisfactory completeness of our hypotheses. For example, we find that the advertising ID and GAID are interchangeably used in privacy policy documents to refer to the Google Ad ID. In that case, we formulate hypotheses for each synonymous term. Besides that, we also tailor the verbal phrase for different objects in forming the hypotheses. For each type of privacy data (including its synonyms), we select all applicable verbal phases from a predefined set of verbs including “collect”, “read”, “request”, etc.
@@
Our second challenge is the complexity of the context in SDK privacy policies. Due to the adopted NLI model is not domain-specifically trained, it is hard to know whether a sentence from the privacy policy (premise) describes the SDK’s behaviors. For example, during our preliminary study, we found a common sentence format from many SDKs’ privacy policies, e.g., “According to GDPR, all collection of device identifiers must be declared in advance”. Conventional inference task tends to determine that sentence implies the collection of device identifiers with a high confidence score. However, it is irrelevant to the actual behaviors of the analyzed SDK. To avoid the disturbance of irrelevant context, we further tailor our hypotheses for each data type. Instead of a single hypothesis to determine if the premise mentions the collection of certain data, we devise three hypotheses corresponding to three different sentiments to regulate the contextual scope, namely positive hypothesis (shown as PH in Figure 2, e.g., “It will collect …”), negative hypothesis (NH, e.g., “It does not collect …”), and irrelevant hypothesis (IR, e.g., “It does not mention whether to collect …”). As a result, we have composed a list of 507 hypotheses covering the 41 pre-identified data types. The full list of hypotheses is provided in Appendix A.2.
[…]
@@
We affirm a data type is requested by an SDK if the NLI model finds at least one sentence from the SDK’s privacy policy that satisfies the criteria stating that the involved SDK will collect that data. Since we formulate three hypotheses for the collection behavior of each combination of verbal phrase and object, we only consider a data type claimed to be collected by an SDK if there exists at least one combination of the associated object (e.g., location) and applicable verbal phrase (e.g., request) satisfying the logical condition as follows:
@@
Intuitively, to infer a specific type of data being collected, the entailment prediction score ($\text{Score}_{p}$) is supposed to be greater than 0.5 to ensure it is always higher than the prediction of the neutral and contradiction hypotheses. However, our preliminary experiments observe a large number of false positives by setting the threshold to 0.5. For that reason, we perform a small-scale analysis to tune the threshold value from 0.5 and above until we find an optimal value that minimizes the number of false positive predictions. Specifically, we only consider SDKs developed by Google from the Google Play SDK Index based on our assumption that those SDKs should provide high-quality privacy policies. We gathered 16 SDKs authored by Google, among which 13 SDKs provide privacy policies. Due to shared privacy policies among some SDKs (i.e., SDKs for Google Firebase), we eventually reviewed five distinct privacy policies. We manually analyzed declared privacy data types from the five privacy policies to set a ground truth. We then fine-tuned the threshold value and discovered that the NLI model could achieve perfect prediction with a threshold of 0.73. For that reason, we adopt 0.73 as the optimal threshold in our large-scale privacy policy analysis.
To validate NLI model performance on analyzing privacy policies of all the collected SDKs, three members from our organizations are invited to perform a manual validation independently. Each of them holds expertise in Android app analysis and development, with one member additionally having proficiency in drafting privacy policy documents. We will discuss the evaluation outcomes of our inference in Section 5.1.
4.4. Exploring Privacy Collection and Potential Leakages
We leverage the static taint analysis techniques to explore privacy collection behaviors and potential exfiltration in SDKs. We assume that a valid privacy data exfiltration is constituted of a collection action that queries one or more privacy data from the OS, and a sharing action that shares the obtained privacy data, directly or indirectly, to the public spaces including the Internet, system settings, and the public storage. Thus, we stipulate an SDK conducts privacy collection if there exists a data flow that starts from a privacy data collection action in its implementation, and this privacy collection would result in a data exfiltration if the data flow sinks at a sensitive sharing action.
@@ 4.4. Exploring Privacy Collection and Potential Leakages
After confirming the list of privacy data, we retrieve the API documents to collect all relevant APIs that contribute to a collection action and include them into the source methods for taint analysis. We refer to an existing research (rasthofer2014machine, ) which proposes a technique named SuSi, which manages to identify and classify over 900 API methods that can potentially leak users’ privacy in Android 4.3 and earlier versions. Considering there has been a decade’s evolution of Android OS since the release of SuSi, many APIs are no longer valid because they have become deprecated and then removed from later OS releases. Meanwhile, new hardware and functionalities, e.g., motion sensors and ultra-wideband (UWB), have emerged during the rapid evolution of the Android ecosystem, and accordingly, the scope of privacy data keeps being extended. For those reasons, we first map the APIs identified in (rasthofer2014machine, ) with the 41 types of privacy data. We iterate the developers’ documents of those APIs and filter out those APIs that are no longer valid in Android Open Source Project (AOSP) implementation. We then manually visit the developer’s document and AOSP source code to collect as many APIs that are associated with the new data types as possible. As a result, we managed to collect 741 source methods. [^3] We count the number of methods by unique method signatures in this paper. The distribution of 741 methods corresponding to the 41 types of privacy data can be found in Table 1.
Table 2. A brief statistics of recognized sink methods
@@
Next, we retrieve all methods that perform a sharing action from Android SDKs. Although there are not many new APIs introduced after Android 4.2 that can be treated as sink methods, we find most sink methods studied in earlier literature including SuSi (rasthofer2014machine, ) are no longer valid in the latest Android versions and therefore be excluded in this study. For example, logging is no longer a feasible approach for sharing data among different apps because an app cannot read logs created by another (google2023loginfo, ). Similarly, methods to access and manage SQLite databases are not counted because the database files are protected by the sandbox mechanism of Android and are not visible to other apps (google2022sandbox, ). In the end, we identify 46 methods from 11 classes and then configure them as sink methods. We present brief statistics of the sink method by their containing classes in Table 2.
The majority of sink methods are relevant to data transfer via network interfaces, reserving 37 out of 46 identified method signatures. Besides that, five methods defined in java.io.Writer and one method from java.io.FileOutputStream are recognized as sink methods that can share sensitive data via public storage. The remaining three sink methods are counted because they can be used for modifying the system settings, which is known as an undocumented channel to share data among different parties on the device (meng2023post, ). We note that writing contents into the system settings requires privileged permission and we will discuss it in Section 6.
Adopting Automatic Taint Analysis
We learned that there are some off-the-shelf automatic static taint analysis tools for Android apps. Our preliminary investigation considers three open-sourced tools, namely FlowDroid (arzt2014flowdroid, ), Mariana Trench (mariana_trench, ), and AppShark (appshark, ). We set up these three tools on our experimental PC, feed four best-selling non-game apps from the Google Play App Store [^4] The four apps used for our preliminary investigation are Facebook, Instagram, Snapchat, and TikTok. into these tools, and explore data flows simultaneously containing a privacy collection action and a sharing action. Our observation shows that AppShark performs the best in terms of time efficiency, success rate of analysis, and capability of handling inter-process communication (IPC). For those reasons, we adopt AppShark for the taint analysis of SDKs on a large scale.
5. Evaluation
Our evaluation aims to assess the performance of our proposed NLI model-based privacy policy analysis and explore the privacy data collection and potential leakages in the collected SDKs. After that, we analyze the landscape of privacy preservation at the SDK level by cross-checking the data requested in the privacy policies and the data collected in SDKs’ implementation.
@@ 5. Evaluation
We implement our privacy policy analysis in Python with PyTorch. The NLI model is provided by a Python package named Transformers. [^5] Can be accessed through URL https://huggingface.co/docs/transformers/index Our static taint analysis is implemented in Java and it supports users to specify target apps and customize rules for taint analysis in a configuration file. For each SDK, we implement a blank Android app to embed it and pass the app package to the automatic taint analysis program. Considering that taint analysis may unavoidably produce false positives, manual validation of the tainted traces is performed to ensure the correctness of our further assessment. We will open-source our implementation and relevant artifacts upon paper acceptance.
@@
5.1. RQ1: Privacy Policy Analysis
Table 3. Evaluation outcomes of privacy policy analysis (only data types requested are displayed)
@@ 5.1. RQ1: Privacy Policy Analysis
We perform NLI model-based analysis on the 109 SDK privacy policies and present the analysis outcomes in Table 3. Our analysis manages to identify 214 data requests from privacy policies. Among the 214 data requests, IP address reserves the largest share, contributing to 47 (21.9%) occurrences. network information and location are the remaining data types among the top three, which are requested in 30 (14.0%) and 28 (13.0%) occurrences, respectively. From the perspective of SDKs, the number of data types claimed by an SDK ranges from 0 to 6, with the average number 1.9.
We also conduct a manual validation to evaluate the performance of our proposed NLI model-based analysis in identifying data requested from privacy policies. After an independent manual inspection of our findings in the 109 privacy policies, we validate 187 data requests covering 19 data types, and accordingly treat these 187 requests as the ground truth in evaluating our analysis approach. Overall, the evaluation outcomes show that our approach is capable of performing privacy policy analysis on a large scale, evidenced by a promising precision of 87.4% and F1 score of 89.5% (see Table 3).
We find our NLI model-based analysis achieves perfect precision in detecting data collection for 10 data types despite 27 false positive cases being observed, i.e., wrongly recognized by the NLI model. Our evaluation outcomes show that account information (6), network information (5), and SMS (4) are the top three data types involved in false positives, indicating our NLI model together with the formulated hypotheses cannot well infer the collection of certain data types. The main reason for these false positives may be the misunderstanding of the adopted LLM to the terms we use to describe the involved privacy data, as well as the scope of certain data types. A typical example is the analysis of collecting privacy data relevant to SMS. Our approach is capable of understanding the literal meaning of SMS but fails to differentiate SMS and other messaging approaches such as e-mails. We find that our adopted NLI model mistakenly entails the collection behavior with high confidence when processing sentences like “we may send email to the users…” in some SDKs. We elaborate more on these false positives with real examples in Appendix A.3.
5.2. RQ2: Privacy Collection and Potential Leakages at the SDK level
Our static taint analysis manages to identify 2318 occurrences of privacy data access in their implementation, covering 27 out of the 41 pre-identified data types. 95 out of 158 (60.1%) SDKs are found to read privacy data, and 34 (21.5%) SDKs have at least one trace tainted. Among these access behaviors, 346 (14.9%) accesses are tainted during the analysis, indicating their data flows contain a sharing action that potentially leaks the accessed privacy data to the public spaces. We present the taint analysis outcomes by data types in Table 4. We also performed manual validation to exclude false positive tainted traces, which we will elaborate on with examples in the remainder of this section.
Table 4. A brief statistics of the taint analysis results (only data types called and/or tainted are displayed)

### v2

@@ III-C Phase 1: Collecting Android SDKs
We identify the Google Play SDK Index as the primary source for collecting Android SDKs. Additionally, considering the uniqueness of mainland China’s Android ecosystem where Google Play infrastructure is absent, we use CAICT as a supplementary source. As a result, we download the latest version of 158 distinct SDKs as of October 2022, among which 139 are from the Google Play SDK Index and 19 are released on the CAICT website. These 158 constitute our previously defined $\mathbb{S}$ (Section II-C) for the later assessment. We also revisit the latest versions of the same group of SDKs in October 2023 and conduct an additional assessment, which we will detail in Section IV-D.
We remark that, at the time of this study, both the Google Play SDK Index and CAICT have not enforced strict regulation or code auditing mechanisms like the Google Play App Store. As a result, there is no guarantee that the SDKs from these sources include comprehensive documentation or privacy policies required by regulations such as GDPR.
@@
Next, we implement a crawler with Selenium [34] to retrieve meta information and privacy policy links from the two sources. The meta information includes the SDK name, provider (developer’s identity), Maven IDs, and functional categories. Such information will be jointly used to uniquely identify SDKs during our subsequent analysis. We then collect the privacy policies from the retrieved links. We will use them to extract the list of personal data that the SDK requests.
The retrieval of privacy policies is found more complex than collecting meta information due to the lack of a compulsory standard for drafting them. For SDKs from the Google Play SDK Index, we rely on the links provided in the “data safety section” section if available, and download the corresponding webpages. Similarly, for SDKs from the CAICT website, we download the webpages from the “SDK privacy policy” link on the SDK details page. We use Google Translate to process all non-English text on these webpages.
Due to the limited availability of privacy policies, we only collected policies for 52 SDKs directly from the two platforms. For those do not provide privacy policy URLs, we manually searched the Internet using keywords like “disclaimer”, “policy”, “disclosure”, and “guidance” to look for any available privacy policies drafted by the SDKs’ developers. As a result, we identify privacy policy webpages for another 57 SDKs. In total, we obtained valid privacy policies for 109 SDKs (69.0% of all SDKs). The text from these webpages will be extracted for further analysis.
III-D Phase 2: Identifying Data Requested in Privacy Policies
After collecting the 109 privacy policies, our next step aims to identify the data requested by each SDK from them. In contrast to traditional app privacy policies that are written in the user-oriented language, SDK privacy policies are more developer-oriented and may contain many domain-specific terminology (e.g., ICCID, BSSID) and technical synonyms (e.g., Google Ad ID, which also appears as Advertising ID, or GAID). This uniqueness makes it challenging for existing methods [35, 32, 33, 36, 37, 38] to extract domain-specific entities accurately. To tackle these challenges, we leverage an LLM-based Natural Language Inference (NLI) model to infer the data collection practice from the privacy policies.
Figure 2: The workflow of our privacy policy analysis to find out what an SDK claims to collect
@@ III-D Phase 2: Identifying Data Requested in Privacy Policies
NLI has proven effective in understanding natural language across different domains [39]. In this work, we use the state-of-the-art LLM-based NLI model named roberta-NLI [^1] The full name of the model is roberta-large-snli_mnli_fever_anli _R1_R2_R3-nli. We shortly write as roberta-NLI to save space. to identify requested data in privacy policies. By parsing sentences from these policies as premises, we input them into the model alongside hypotheses based on 41 pre-identified data types (refer to Section III-B), allowing the model to assess the entailment relationship between them. We briefly illustrate our approach in Figure 2.
There are two key challenges arose during the hypothesis formulation. The first challenge is the missing of a canonical writing standard of SDK privacy policies. Privacy policies often use developer-oriented language with technical synonyms, requiring us to manually gather synonyms (e.g., the advertising ID and GAID are interchangeably used to refer to the Google Ad ID) to ensure hypothesis completeness. Besides that, we also tailored hypotheses by matching different verb phases (e.g., “collect”, “read”) to each data type and its synonyms to maximize the entailment score when the collection of the inferred data type is mentioned in the privacy policy.
@@
The second challenge is handling the complex context of SDK privacy policies, as the NLI model is not domain-specifically trained. For example, we found common sentence formats in privacy policies, such as “According to GDPR, all collection of device identifiers must be declared in advance.” Conventional inference models might incorrectly conclude with high confidence that this implies the collection of device identifiers, but it is irrelevant to the actual behaviors of the SDK being analyzed. To address this issue, we created three tailored hypotheses for each combination of verbal phrase and object, namely the positive hypothesis (shown as PH in Figure 2, e.g., “It will collect …”), negative hypothesis (NH, e.g., “It does not collect …”), and irrelevant hypothesis (IR, e.g., “It does not mention whether to collect …”). As a result, we have composed a list of 507 hypotheses covering the 41 pre-identified data types.
@@
Given three hypotheses have been defined for each combination of verbal phrase and object, we only consider a data type claimed to be collected by an SDK if there exists at least one combination of the associated object (e.g., location) and applicable verbal phrase (e.g., request) satisfying the logical condition as follows:
@@
Intuitively, to infer a specific type of data being collected, the entailment prediction score ($\text{Score}_{p}$) is supposed to be greater than 0.5 to ensure it is always higher than the prediction of the neutral and contradiction hypotheses. However, our preliminary experiments observe a large number of false positives by setting the threshold to 0.5. For that reason, we perform a small-scale analysis to tune the threshold value from 0.5 and above until we find an optimal value that minimizes the number of false positive predictions. Specifically, we assume that Google should provide high-quality privacy policies for their own SDKs, and therefore we collected five distinct privacy policy shared by 13 Google in-house SDKs. We manually analyzed declared privacy data types from the five privacy policies to set a ground truth. We then fine-tuned the threshold value and discovered that the NLI model could achieve perfect prediction with a threshold of 0.73. For that reason, we adopt 0.73 as the optimal threshold in our large-scale privacy policy analysis.
III-E Phase 3: Exploring Privacy Collection and Leakages
We leverage the static taint analysis techniques to explore privacy collection behaviors and potential exfiltration in SDKs. We define a valid privacy data exfiltration is constituted of a collection action that queries one or more privacy data from the OS, and a sharing action that shares the obtained privacy data, directly or indirectly, to the public spaces including the Internet, system settings, and the public storage. Thus, we consider that an SDK engages in privacy collection if its implementation contains a data flow originating from a privacy data collection action. This privacy collection leads to exfiltration if the data flow ends at a sensitive sharing action.
@@ III-E Phase 3: Exploring Privacy Collection and Leakages
We aim to collect all relevant APIs that contribute to a collection action and include them into the source methods for taint analysis. To this end, We refer to an existing research [40] which proposes a technique named SuSi, which classifies over 900 API methods that can potentially leak users’ privacy in Android 4.3 and earlier versions. Considering there have been multiple version’s evolutions since Android 4.3 that the API rules may have experienced significant change, we first map the APIs identified in [40] with the 41 types of privacy data. We iterate the developers’ documents of those APIs and filter out those APIs that are no longer valid in Android Open Source Project (AOSP) implementation. We then manually visit the developer’s document and AOSP source code to collect as many APIs that are associated with the new data types as possible. As a result, we collect 741 source methods. [^2] We count the number of methods by unique method signatures in this paper. The distribution of 741 methods corresponding to the 41 types of privacy data can be found in Table I.
TABLE II: A brief statistics of recognized sink methods
@@
Next, we retrieve all methods that perform a sharing action from Android SDKs. Although there are not many new APIs introduced after Android 4.2 that can be treated as sink methods, we find most sink methods studied in earlier literature including SuSi [40], such as logging and SQLite databases, are no longer valid in the latest Android versions and therefore be excluded in this study [41, 42]. In the end, we identify 46 methods from 11 classes and then configure them as sink methods. We present a brief statistics of the sink methods by their containing classes in Table II.
The majority of sink methods are relevant to data transfer via network interfaces, reserving 37 out of 46 identified method signatures. Besides that, six method are recognized because they can share sensitive data via public storage. The remaining three sink methods are counted because they can be used for modifying the system settings, which is known as an undocumented channel to share data among different parties on the device [43]. We note that writing contents into the system settings requires privileged permission and we will discuss it in Section V.
IV Evaluation
Our evaluation aims to assess the performance of our proposed NLI model-based privacy policy analysis and explore the privacy data collection and potential leakages in the SDKs. After that, we analyze the privacy preservation landscape at the SDK level by cross-checking the data requested in the privacy policies and the data collected in SDKs’ implementation (refer to Phase 4 of Figure 1).
@@ IV Evaluation
We implement our privacy policy analysis in Python with PyTorch. The NLI model is provided by a Python package named Transformers. [^3] Available at https://huggingface.co/docs/transformers/index
Adoption of Automatic Taint Analysis
We learned that there are some off-the-shelf automatic static taint analysis tools for Android apps. Our preliminary investigation considers three open-sourced tools, namely FlowDroid [44], Mariana Trench [45], and AppShark [46]. We set up these three tools on our experimental PC, feed four best-selling non-game apps from the Google Play App Store, namely Facebook, Instagram, Snapchat, and TikTok, into these tools, and explore data flows simultaneously containing a privacy collection action and a sharing action. Our observation shows that AppShark performs the best in terms of time efficiency, success rate of analysis, and capability of handling inter-process communication (IPC). For those reasons, we adopt AppShark for the taint analysis of SDKs on a large scale. For each SDK, we implement a blank Android app to embed it and pass the app package to the automatic taint analysis program.
@@
IV-A RQ1: Privacy Policy Analysis Performance
TABLE III: Evaluation outcomes of privacy policy analysis (only data types requested are displayed)
@@ IV-A RQ1: Privacy Policy Analysis Performance
We perform NLI model-based analysis on the 109 SDK privacy policies and present the analysis outcomes in Table III. Our analysis identifies 214 data requests from privacy policies. Among the 214 data requests, IP address reserves the largest share, contributing to 47 (21.9%) occurrences. network information and location are the remaining data types among the top three. From the perspective of SDKs, the number of data types claimed by an SDK ranges from 0 to 6, with the average number 1.9.
We then conduct a manual inspection of our findings in the 109 privacy policies. As a result, we find 187 verified data requests covering 19 data types, and accordingly treat these requests as the ground truth in evaluating our analysis approach. The results demonstrate that our approach can effectively analyze privacy policies on a large scale, achieving a promising precision of 87.4% and an F1 score of 89.5%. (see Table III).
We find our NLI model-based analysis achieves perfect precision in detecting data collection for 10 data types despite 27 false positive cases being observed, i.e., wrongly recognized by the NLI model. Our evaluation outcomes show that account information (6), network information (5), and SMS (4) are the top three data types involved in false positives, indicating the model struggled to infer data collection for certain types. These errors likely stem from the model’s misinterpretation of terms and data type scopes. For example, although our NLI model correctly recognizes SMS, it fails to differentiate SMS and other messaging methods like emails, leading to high-confidence false positives while dealing with sentences like “we may send email to the users…” as SMS collection.
IV-B RQ2: Privacy Collection and Potential Leakages at the SDK Level
Our static taint analysis identified 2318 instances of privacy data access across 27 of the 41 pre-identified data types. Out of 158 SDKs, 95 (60.1%) were found to read privacy data, with 34 (21.5%) having at least one tainted trace. The analysis revealed that 346 (14.9%) of these accesses were tainted, suggesting potential privacy leakage into public spaces. We present the taint analysis outcomes by data types in Table IV.
TABLE IV: A brief statistics of the taint analysis results (only data types called and/or tainted are displayed)